ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: 65 Liandrat, J., IRPHE/ESM2, La jetee, Technopole de Chateau Gombert, Marseille, France Wavelets, from theory to efficients numerical algorithms In this lecture we will review some links between theoretical results on wavelet approximation for functions and operators, and recently developped adaptive algorithms for PDE resolution. We will show that a good understanding of the properties of wavelet families lead to really efficient numerical algorithms. The results of some numerical tests will be shown.